# ICPO: Intrinsic Confidence-Driven Group Relative Preference Optimization for Efficient Reinforcement Learning

### v3

@@ C.1 Aspect 1: Calibrate Policy Updates
Figure 6: Scatter plot of the log-probability distributions of responses within a batch, where each batch contains 768 groups and each group comprises 5 responses.
@@
We randomly selected a training batch during the overall training process and statistically analyzed the log-probability distributions of different responses within the group. Furthermore, we employed two sets of real in-group response examples to illustrate the aforementioned theorem. Figure 6 presents the log-probability distributions of responses within a training batch. It can be observed that the log-probability of most responses fluctuate between -0.15 and 0.0, while a small number of extreme responses exhibit log-probability reaching -0.25.

### v4

@@ C.1 Aspect 1: Calibrate Policy Updates
Figure 6: Scatter plot of responses log-probability distributions within a batch, where each batch contains 768 groups and each group comprises 5 responses.
@@
We randomly selected a training batch during the overall training process and statistically analyzed the log-probability distributions of different responses within the group. Furthermore, we employed two sets of real in-group response examples to illustrate the aforementioned theorem. Figure 6 presents the log-probability distributions of responses within a training batch. It can be observed that the log-probability of most responses fluctuate between -2.0 and 0.0, while a small number of extreme responses exhibit log-probability reaching -3.0.
@@
Log-probability: $P_{1}(-0.135)$, $P_{2}(-0.583)$, $P_{3}(-0.216)$, $P_{4}(-0.050)$, $P_{5}(-0.407)$
- $S^{p}_{2}=0.4\cdot(\frac{-0.407}{-0.583}+\frac{-0.216}{-0.583}+\frac{-0.135}{-0.583}+\frac{-0.050}{-0.583})-0.4\cdot(-0.050)=0.5744$
- $S^{p}_{5}=0.4\cdot(\frac{-0.216}{-0.407}+\frac{-0.135}{-0.407}+\frac{-0.050}{-0.407})-0.4\cdot(-0.050)=0.4141$
- $S^{p}_{3}=0.4\cdot(\frac{-0.135}{-0.216}+\frac{-0.050}{-0.216})-0.4\cdot(-0.050)=0.3626$
- $S^{p}_{1}=0.4\cdot(\frac{-0.050}{-0.135})-0.4\cdot(-0.050)=0.1681$
- $S^{p}_{4}=-0.4\cdot(-0.050)=0.020$
- $R^{final}_{2}=0.1000+1.0\cdot\operatorname{min}(0.5744,\frac{0.1000}{2.0})=0.15$
- $R^{final}_{5}=1.0000+1.0\cdot\operatorname{min}(0.4141,\frac{1.0000}{2.0})=1.4141$
- $R^{final}_{3}=1.0000+1.0\cdot\operatorname{min}(0.3626,\frac{1.0000}{2.0})=1.3626$
- $R^{final}_{1}=0.1000+1.0\cdot\operatorname{min}(0.1681,\frac{0.1000}{2.0})=0.15$
- $R^{final}_{4}=0.1000+1.0\cdot\operatorname{min}(0.020,\frac{0.1000}{2.0})=0.120$
- ICPO: $A_{1}(-0.7996)$, $A_{2}(-0.7996)$, $A_{3}(1.1819)$, $A_{4}(-0.8986)$, $A_{5}(1.2660)$
Log-probability: $P_{1}(-0.286)$, $P_{2}(-0.287)$, $P_{3}(-0.144)$, $P_{4}(-0.066)$, $P_{5}(-2.356)$
- $S^{p}_{5}=0.4\cdot(\frac{-0.287}{-2.356}+\frac{-0.286}{-2.356}+\frac{-0.144}{-2.356}+\frac{-0.066}{-2.356})-0.4\cdot(-0.066)=0.1593$
- $S^{p}_{2}=0.4\cdot(\frac{-0.286}{-0.287}+\frac{-0.144}{-0.287}+\frac{-0.066}{-0.287})-0.4\cdot(-0.066)=0.7177$
- $S^{p}_{1}=0.4\cdot(\frac{-0.144}{-0.286}+\frac{-0.066}{-0.286})-0.4\cdot(-0.066)=0.3201$
- $S^{p}_{3}=0.4\cdot(\frac{-0.066}{-0.144})-0.4\cdot(-0.066)=0.2097$
- $S^{p}_{4}=-0.4\cdot(-0.066)=0.0264$
- $R^{final}_{5}=1.0000+1.0\cdot\operatorname{min}(0.1593,\frac{1.0000}{2.0})=1.1593$
- $R^{final}_{2}=0.1000+1.0\cdot\operatorname{min}(0.7177,\frac{0.1000}{2.0})=0.15$
- $R^{final}_{1}=0.1000+1.0\cdot\operatorname{min}(0.3201,\frac{0.1000}{2.0})=0.15$
- $R^{final}_{3}=1.0000+1.0\cdot\operatorname{min}(0.2097,\frac{1.0000}{2.0})=1.2097$
- $R^{final}_{4}=0.1000+1.0\cdot\operatorname{min}(0.0264,\frac{0.1000}{2.0})=0.1264$
- ICPO: $A_{1}(-0.8006)$, $A_{2}(-0.8006)$, $A_{3}(1.2733)$, $A_{4}(-0.8468)$, $A_{5}(1.1746)$
Log-probability: $P_{1}(-0.135)$, $P_{2}(-0.583)$, $P_{3}(-0.216)$, $P_{4}(-0.050)$, $P_{5}(-0.407)$
- $S^{p}_{2}=0.4\cdot(\frac{-0.407}{-0.583}+\frac{-0.216}{-0.583}+\frac{-0.135}{-0.583}+\frac{-0.050}{-0.583})-0.4\cdot(-0.050)=0.5744$
- $S^{p}_{5}=0.4\cdot(\frac{-0.216}{-0.407}+\frac{-0.135}{-0.407}+\frac{-0.050}{-0.407})-0.4\cdot(-0.050)=0.4141$
- $S^{p}_{3}=0.4\cdot(\frac{-0.135}{-0.216}+\frac{-0.050}{-0.216})-0.4\cdot(-0.050)=0.3626$
- $S^{p}_{1}=0.4\cdot(\frac{-0.050}{-0.135})-0.4\cdot(-0.050)=0.1681$
- $S^{p}_{4}=-0.4\cdot(-0.050)=0.020$
- $R^{final}_{2}=0.1000+1.0\cdot\operatorname{min}(0.5744,\frac{0.1000}{2.0})=0.15$
- $R^{final}_{5}=0.7000+1.0\cdot\operatorname{min}(0.4141,\frac{0.7000}{2.0})=1.05$
- $R^{final}_{3}=1.0000+1.0\cdot\operatorname{min}(0.3626,\frac{1.0000}{2.0})=1.3626$
- $R^{final}_{1}=0.4000+1.0\cdot\operatorname{min}(0.1681,\frac{0.4000}{2.0})=0.5681$
- $R^{final}_{4}=0.4000+1.0\cdot\operatorname{min}(0.020,\frac{0.4000}{2.0})=0.420$
- ICPO $w$ $R^{Noisy}$: $A_{1}(-0.3243)$, $A_{2}(-1.2788)$, $A_{3}(1.4900)$, $A_{4}(-0.6624)$, $A_{5}(0.7759)$